# Tailoring Adversarial Attacks on Deep Neural Networks for Targeted Class Manipulation Using DeepFool Algorithm

### v4

@@ Abstract
The susceptibility of deep neural networks (DNNs) to adversarial attacks undermines their reliability across numerous applications, underscoring the necessity for an in-depth exploration of these vulnerabilities and the formulation of robust defense strategies. The DeepFool algorithm by Moosavi-Dezfooli et al. (2016) represents a pivotal step in identifying minimal perturbations required to induce misclassification of input images. Nonetheless, its generic methodology falls short in scenarios necessitating targeted interventions. Additionally, previous research studies have predominantly concentrated on the success rate of attacks without adequately addressing the consequential distortion of images, the maintenance of image quality, or the confidence threshold required for misclassification. To bridge these gaps, we introduce the Enhanced Targeted DeepFool (ET DeepFool) algorithm, an evolution of DeepFool that not only facilitates the specification of desired misclassification targets but also incorporates a configurable minimum confidence score. Our empirical investigations demonstrate the superiority of this refined approach in maintaining the integrity of images and minimizing perturbations across a variety of DNN architectures. Unlike previous iterations, such as the Targeted DeepFool by Gajjar et al. (2022), our method grants unparalleled control over the perturbation process, enabling precise manipulation of model responses. Preliminary outcomes reveal that certain models, including AlexNet and the advanced Vision Transformer, display commendable robustness to such manipulations. This discovery of varying levels of model robustness, as unveiled through our confidence level adjustments, could have far-reaching implications for the field of image recognition. Our code will be made public upon acceptance of the paper.
@@ 3.1 Vanilla DeepFool
In a multi-class classification setting, each class can be thought of as being separated by a boundary, known as a hyperplane. When an input, represented as $\bm{x}$, is classified, it gets assigned to the class on the side of the boundary where it falls. The DeepFool algorithm works by finding the nearest boundary and slightly adjusting $\bm{x}$ so that it crosses this boundary and gets misclassified. This process is repeated until the input is misclassified, and the algorithm returns the total amount of changes (called perturbations), denoted as $\bm{\hat{r}}$, needed to cause the misclassification. To find the closest boundary, denoted as $\hat{l}$, the algorithm first calculates a vector $\bm{w}^{\prime}_{k}$ that points toward the boundary between the current predicted class and another class $k$. This vector is found by subtracting the gradient of the current predicted class from the gradient of class $k$:
@@ 3.2 Targeted DeepFool
Basic Algorithm. This is an iterative approach which works by making small changes or perturbations to an image. The goal of these perturbations is to gradually alter the image so that a classifier starts to see it as belonging to a different, specific target class rather than its original class. The process is iterative: in each step, the algorithm computes the direction and magnitude of the perturbation that will push the image closer to being misclassified as the target class. This is done by moving the image towards the hyperplane of the classifier that separates the original class from the target class. The algorithm continues this process, until one of two things happens: either the classifier incorrectly labels the image as the target class, or the amount of perturbation applied to the image reaches a pre-set limit, beyond which further changes might make the image too distorted.
Recursive Algorithm. This approach extends the basic idea by addressing situations where the basic algorithm might get “stuck” on an intermediate step, unable to push the image all the way to the target class. This can happen because the decision boundaries in a deep neural network are often non-linear, making it difficult for a simple iterative process to find a path to the target class. In the recursive approach, if the basic algorithm fails to reach the target classification after a certain number of iterations, the algorithm resets and starts the process again, but this time from the current altered state of the image. This refines the image’s perturbations in each pass, which helps the image navigate around difficult regions of the decision boundaries. This process continues until the image is successfully misclassified as the target class or a maximum number of recursion attempts is reached. This recursive strategy makes the approach more robust and often more effective than the basic iterative method.
@@ 4.1 Enhanced Targeted DeepFool
Algorithm 1 DeepFool: multi-class case
1:Input: Image $\bm{x}$, classifier $f$.
@@
3:Output: Perturbation $\bm{\hat{r}}$.
4:Initialize $\bm{x}_{0}\leftarrow\bm{x},i\leftarrow 0$.
@@
Algorithm 2 Proposed: Enhanced Targeted DeepFool
1:Input: Image $\bm{x}$, classifier $f$.
2:target class $t$, minimum confidence $\bm{c}_{min}$.
3:Output: Perturbation $\bm{\hat{r}}$.
@@
Comparison between the original DeepFool and our proposed Enhanced Targeted DeepFool algorithm
To start with, we run the while loop until the current label is not equal to the target label instead of running it until the image is misclassified. Additionally, we have eliminated the inner for loop, as shown in line 6 of Algorithm 1. This loop calculates the gradients for all classes except the original one, which is not necessary when we focus on a single target class. The removal of this loop significantly reduces the computational complexity of the algorithm, effectively decreasing it by a factor of $O(C)$, where $C$ is the number of classes.
@@ 5.1 Dataset
We use six different datasets; MNIST [38], CIFAR-10 [39], EuroSAT [40], GTSRB [41], HAM10000 [42] and the ILSVRC2012 [43] also known as the ImageNet dataset for our experiments. The MNIST dataset consists of a collection of $28\times 28$ grayscale images of handwritten digits (0 to 9), along with their corresponding labels. It contains 60,000 training images and 10,000 test images. The CIFAR-10 dataset is a collection of color images, each $32\times 32$ pixels in size, spread across 10 different classes. These classes include common objects such as airplanes, automobiles, birds, cats, deer, dogs, frogs, horses, ships, and trucks with 50000 training images and 10000 test images. The EuroSAT dataset is a benchmark dataset for satellite image classification, derived from the Sentinel-2 satellite imagery. It contains 27,000 labeled $64\times 64$ images across 10 classes, including various land cover types like residential areas, forests, and water bodies. The German Traffic Sign Recognition Benchmark (GTSRB) is a dataset used for training and evaluating traffic sign recognition systems. It consists of over 50,000 images of 43 different traffic sign classes, collected under various lighting and weather conditions. The dataset is commonly used for autonomous driving applications. The HAM10000 (Human Against Machine with 10000 training images) dataset comprises 10,015 dermatoscopic images collected from diverse populations and imaging modalities. These images are categorized into various skin lesion types, including melanoma, nevus, and benign keratosis. The ImageNet dataset contains 50,000 images with a thousand different classes. It is widely regarded as a benchmark dataset in the field of computer vision and has played a crucial role in advancing the development of deep learning models. The dataset is large and diverse which makes it a comprehensive representation of real-world visual data. Due to its size and diversity, the models pre-trained on this dataset can learn rich feature representation that captures a good amount of visual information.
@@ 7 Discussion
Table 8: Comparison among existing attacks and our proposed attack on the ImageNet dataset. Here, * represents a variable that can range from 0 to 100, depending on our preference
@@
Table 9: Comparison among existing attacks and our proposed attack on the MNIST dataset. Here, * represents a variable that can range from 0 to 100, depending on our preference
@@
Table 10: Comparison among existing attacks and our proposed attack on the CIFAR-10 dataset. Here, * represents a variable that can range from 0 to 100, depending on our preference
@@
The Targeted DeepFool algorithm by Gajjar et al. uses the MNIST dataset, which is a $28\times 28$ pixel grayscale dataset of numbers, and the CIFAR-10 dataset which only contains 10 classes [18]. The images in these datasets are not suitable to compare against real-life images. Their approach has a 97.83% success rate which is a bit higher than our 96.65%. However, our approach generates images that have a lower perturbation rate (1.29%) as shown in Table 9. In Table 10, we can see that they get a 77% success rate for CIFAR-10 whereas our approach gets a 94.22% success rate with significantly less perturbation rate (3.04%) than their 11%. They also claim that the adversarial images are visually indistinguishable from the original ones, however, this claim is questionable, as the images are too small to assess the level of distortion. Furthermore, although their basic iterative approach shares the same time complexity as ours, it under performs compared to their most efficient method, the recursive algorithm, which is more time-consuming, as demonstrated in Table 13. Additionally, in contrast to the majority of untargeted attacks in the literature, our study does not seek to elucidate the degradation in model performance when retrained with perturbed images; consequently, we have chosen not to incorporate this aspect in our analysis.
@@ 9 Acknowledgment
We would like to acknowledge BRAC University and the University of Alabama at Birmingham for providing computational resources, which were essential for the completion of this research.

### v5

@@ Abstract
The susceptibility of deep neural networks (DNNs) to adversarial attacks undermines their reliability across numerous applications, underscoring the necessity for an in-depth exploration of these vulnerabilities and the formulation of robust defense strategies. The DeepFool algorithm by Moosavi-Dezfooli et al. (2016) represents a pivotal step in identifying minimal perturbations required to induce misclassification of input images. Nonetheless, its generic methodology falls short in scenarios necessitating targeted interventions. Additionally, previous research studies have predominantly concentrated on the success rate of attacks without adequately addressing the consequential distortion of images, the maintenance of image quality, or the confidence threshold required for misclassification. To bridge these gaps, we introduce the Enhanced Targeted DeepFool (ET DeepFool) algorithm, an evolution of DeepFool that not only facilitates the specification of desired misclassification targets but also incorporates a configurable minimum confidence score. Our empirical investigations demonstrate the superiority of this refined approach in maintaining the integrity of images and minimizing perturbations across a variety of DNN architectures. Unlike previous iterations, such as the Targeted DeepFool by Gajjar et al. (2022), our method grants unparalleled control over the perturbation process, enabling precise manipulation of model responses. Preliminary outcomes reveal that certain models, including AlexNet and the advanced Vision Transformer, display commendable robustness to such manipulations. This discovery of varying levels of model robustness, as unveiled through our confidence level adjustments, could have far-reaching implications for the field of image recognition. Our code is available at https://github.com/FazleLabib/et_deepfool.
@@ 3.1 Vanilla DeepFool
In a multi-class classification setting as seen on Algorithm 1, each class can be thought of as being separated by a boundary, known as a hyperplane. When an input, represented as $\bm{x}$, is classified, it gets assigned to the class on the side of the boundary where it falls. The DeepFool algorithm works by finding the nearest boundary and slightly adjusting $\bm{x}$ so that it crosses this boundary and gets misclassified. This process is repeated until the input is misclassified, and the algorithm returns the total amount of changes (called perturbations), denoted as $\bm{\hat{r}}$, needed to cause the misclassification. To find the closest boundary, denoted as $\hat{l}$, the algorithm first calculates a vector $\bm{w}^{\prime}_{k}$ that points toward the boundary between the current predicted class and another class $k$. This vector is found by subtracting the gradient of the current predicted class from the gradient of class $k$:
@@ 3.2 Targeted DeepFool
Basic Targeted DeepFool Algorithm: This is an iterative approach as seen on Algorithm 2, which works by making small changes or perturbations to an image. The goal of these perturbations is to gradually alter the image so that a classifier starts to see it as belonging to a different, specific target class rather than its original class. The process is iterative: in each step, the algorithm computes the direction and magnitude of the perturbation that will push the image closer to being misclassified as the target class. This is done by moving the image towards the hyperplane of the classifier that separates the original class from the target class. The algorithm continues this process, until one of two things happens: either the classifier incorrectly labels the image as the target class, or the amount of perturbation applied to the image reaches a pre-set limit, beyond which further changes might make the image too distorted.
Recursive Targeted DeepFool Algorithm: This approach extends the basic idea by addressing situations where the basic algorithm might get “stuck” on an intermediate step, unable to push the image all the way to the target class. This can happen because the decision boundaries in a deep neural network are often non-linear, making it difficult for a simple iterative process to find a path to the target class. In the recursive approach, if the basic algorithm fails to reach the target classification after a certain number of iterations, the algorithm resets and starts the process again, but this time from the current altered state of the image. This refines the image’s perturbations in each pass, which helps the image navigate around difficult regions of the decision boundaries. This process continues until the image is successfully misclassified as the target class or a maximum number of recursion attempts is reached. This recursive strategy makes the approach more robust and often more effective than the basic iterative method.
@@ 4.1 Enhanced Targeted DeepFool
Algorithm 1
     DeepFool: multi-class case
1:Input: Image $\bm{x}$, classifier $f$
@@
3:Output: Perturbation $\bm{\hat{r}}$
4:Initialize $\bm{x}_{0}\leftarrow\bm{x},i\leftarrow 0$
@@
Algorithm 2
     Basic Targeted DeepFool
1:Input: Image $\bm{x}$, classifier $h$,
2:     target class $t$, total perturbation $p_{tot}$
3:Output: Total perturbation $p_{tot}$
4:Initialize $\bm{x}_{0}\leftarrow\bm{x}$, $i\leftarrow 0$
5:Set threshold (maximum perturbation)
6:while $\arg\max_{j}h_{j}(\bm{x}_{i})\neq t$
7:                  or $p_{i}<\text{threshold}$ do
8:
9:     $p_{i}\leftarrow\frac{|h_{t}(\bm{x}_{i})-h_{j}(\bm{x}_{i})({x}_{i})|}{||g_{t}-g_{j}(\bm{x}_{i})||_{2}^{2}}(g_{t}-g_{j}(\bm{x}_{i}))$
10:
11:     $\bm{x}_{i+1}\leftarrow\bm{x}_{i}+p_{i}$
12:     $i\leftarrow i+1$
13:end while
14:$p_{tot}\leftarrow\sum_{i}p_{i}$
15:
16:return $p_{tot}$
Algorithm 3 Proposed:
     Enhanced Targeted DeepFool
1:Input: Image $\bm{x}$, classifier $f$
2:  target class $t$, min confidence $\bm{c}_{min}$
3:Output: Perturbation $\bm{\hat{r}}$
@@
Comparison among the original DeepFool, basic Targeted DeepFool, and our proposed Enhanced Targeted DeepFool algorithms
To start with, we run the while loop until the current label is not equal to the target label instead of running it until the image is misclassified. Additionally, we have eliminated the inner for loop, as compared to line 6 of Algorithm 1. This loop calculates the gradients for all classes except the original one, which is not necessary in the proposed case when we focus on a single target class. In Algorithm 1 there is a for loop inside a while loop. The while loop is responsible for the complexity as $O(N)$ and the inner for loop is responsible for the complexity as $O(C)$, both create the complexity as $O(NC)$ for Algorithm 1. Here $N$ is the maximum number of iterations and $C$ is the number of classes. On the other hand, in our proposed method (Algorithm 3), we remove the inner for loop from the inside of the while loop as compared to Algorithm 1. Therefore, the complexity of the proposed method depends only upon the while loop and its $O(N)$. The removal of the inner for loop from inside the while loop significantly reduces the computational complexity of Algorithm 3, effectively decreasing the complexity by a factor of $O(C)$, where $C$ is the number of classes.
@@ 5.1 Dataset
We use six different datasets; MNIST [38], CIFAR-10 [39], EuroSAT [40], GTSRB [41], HAM10000 [42] and the ILSVRC2012 [43] also known as the ImageNet dataset for our experiments. The MNIST dataset consists of a collection of $28\times 28$ grayscale images of handwritten digits (0 to 9), along with their corresponding labels. It contains 60,000 training images and 10,000 test images. The CIFAR-10 dataset is a collection of color images, each $32\times 32$ pixels in size, spread across 10 different classes. These classes include common objects such as airplanes, automobiles, birds, cats, deer, dogs, frogs, horses, ships, and trucks with 50000 training images and 10000 test images. The EuroSAT dataset is a benchmark dataset for satellite image classification, derived from the Sentinel-2 satellite imagery. It contains 27,000 labeled $64\times 64$ images across 10 classes, including various land cover types like residential areas, forests, and water bodies. The German Traffic Sign Recognition Benchmark (GTSRB) is a dataset used for training and evaluating traffic sign recognition systems. It consists of over 50,000 images of 43 different traffic sign classes, collected under various lighting and weather conditions. The dataset is commonly used for autonomous driving applications. The HAM10000 (Human Against Machine with 10000 training images) dataset comprises 10,015 dermatoscopic images collected from diverse populations and imaging modalities. These images are categorized into various skin lesion types, including melanoma, nevus, and benign keratosis. The ImageNet dataset contains 1,281,167 training images, 50,000 validation images and 100,000 test images with a thousand different classes. It is widely regarded as a benchmark dataset in the field of computer vision and has played a crucial role in advancing the development of deep learning models. The dataset is large and diverse which makes it a comprehensive representation of real-world visual data. Due to its size and diversity, the models pre-trained on this dataset can learn rich feature representation that captures a good amount of visual information.
For our experiments, we used the test sets provided with MNIST, CIFAR-10, EuroSAT, GTSRB, and HAM10000. On the other hand, as the label of the test set of the ImageNet dataset is not public, therefore, we used the validation dataset for the experiment of our study. It is worth mentioning that, the validation images of the ImageNet dataset are used in several research studies in the literature [44, 45, 46, 47].
@@ 7 Discussion
Table 8: Comparison among existing attacks and our proposed attack on the ImageNet dataset. Reported values are sourced from their respective papers. Fields marked with ’-’ were not reported in the original studies. ’*’ represents a variable that can range from 0 to 100, depending on our preference
@@
Table 9: Comparison among existing attacks and our proposed attack on the MNIST dataset. Reported values are sourced from their respective papers. Fields marked with ’-’ were not reported in the original studies. ’*’ represents a variable that can range from 0 to 100, depending on our preference
@@
Table 10: Comparison among existing attacks and our proposed attack on the CIFAR-10 dataset. Reported values are sourced from their respective papers. Fields marked with ’-’ were not reported in the original studies. ’*’ represents a variable that can range from 0 to 100, depending on our preference
@@
The Targeted DeepFool algorithm by Gajjar et al., uses the MNIST dataset, which is a $28\times 28$ pixel grayscale dataset of numbers, and the CIFAR-10 dataset which only contains 10 classes [18]. The images in these datasets are not suitable to compare against real-life images. Their approach has a 97.83% success rate which is a bit higher than our 96.65%. However, our approach generates images that have a lower perturbation rate (1.29%) as shown in Table 9. In Table 10, we can see that they get a 77% success rate for CIFAR-10 whereas our approach gets a 94.22% success rate with significantly less perturbation rate (3.04%) than their 11%. They also claim that the adversarial images are visually indistinguishable from the original ones, however, this claim is questionable, as the images are too small to assess the level of distortion. Furthermore, although their basic iterative approach shares the same time complexity as ours, it under performs compared to their most efficient method, the recursive algorithm, which is more time-consuming, as demonstrated in Table 13. Additionally, in contrast to the majority of untargeted attacks in the literature, our study does not seek to elucidate the degradation in model performance when retrained with perturbed images; consequently, we have chosen not to incorporate this aspect in our analysis.
@@ 9 Acknowledgment
We gratefully acknowledge the computational resources provided by BRAC University and also the University of Alabama at Birmingham IT-Research Computing group for high-performance computing (HPC) support and CPU time on the Cheaha compute cluster, which was essential for the completion of this research.